### 27 Mar 2026 v1

# DFM-VLA: Iterative Action Refinement for Robot Manipulation via Discrete Flow Matching

Abstract
Vision–Language–Action (VLA) models that encode actions using a discrete tokenization scheme are increasingly adopted for robotic manipulation, but existing decoding paradigms remain fundamentally limited. Whether actions are decoded sequentially by autoregressive VLAs or in parallel by discrete diffusion VLAs, once a token is generated, it is typically fixed and cannot be revised in subsequent iterations, so early token errors cannot be effectively corrected later. We propose DFM-VLA, a discrete flow matching VLA for iterative refinement of action tokens. DFM-VLA models a token-level probability velocity field that dynamically updates the full action sequence across refinement iterations. We investigate two ways to construct the velocity field: an auxiliary velocity-head formulation and an action-embedding-guided formulation. Our framework further adopts a two-stage decoding strategy with an iterative refinement stage followed by deterministic validation for stable convergence. Extensive experiments on CALVIN, LIBERO, and real-world manipulation tasks show that DFM-VLA consistently outperforms strong autoregressive, discrete diffusion, and continuous diffusion baselines in manipulation performance while retaining high inference efficiency. In particular, DFM-VLA achieves an average success length of 4.44 on CALVIN and an average success rate of 95.7% on LIBERO, highlighting the value of action refinement via discrete flow matching for robotic manipulation. Our project is available https://chris1220313648.github.io/DFM-VLA/
[…]
Jiayi Chen at
1 Introduction
Figure 1: Comparison of decoding paradigms. 1) Autoregressive decoding requires as many steps as the action sequence length, while 2) Discrete Diffusion enables faster generation through parallel token updates. However, both have the same limitation that once an erroneous token is produced, it cannot be corrected in later iterations. We refer to this phenomenon as irreversible commitment. In contrast, 3) our DFM-VLA performs full-sequence action refinement at every iteration, allowing token-level correction and improving action quality for robotic manipulation.
Vision–Language–Action (VLA) (univla; Pi0; gr00t; cui2025openhelix; kim2024openvla; song2025reconvla; wang2025vlaadapter; zhong2026dualcot; yan2026svam) models have become a promising foundation for robotic manipulation, where policies map language instructions and visual observations to executable actions. A common and effective design is to discretize actions into tokens, which enables scalable training with Vision-Language Models (VLMs) backbones and leverages their strong capabilities in vision and language understanding.
Most existing discrete VLA systems fall into two families. Autoregressive (AR) methods (brohan2022rt; kim2024openvla; univla) decode tokens sequentially with a next-token prediction objective. However, their inherently left-to-right decoding order makes it difficult to revise erroneous tokens once they are emitted. Discrete diffusion (DD) methods (song2025accelerating; liang2025discrete; wen2025llada) improve parallelism and often reduce latency, but they may still produce erroneous tokens in early iterations under confidence-guided decoding (see Figure˜1). As a result, AR and DD paradigms struggle with a shared issue in robot control: early decoding errors cannot be effectively corrected and therefore propagate through the action chunk, degrading downstream robotic task performance. Recent Discrete Flow Matching (DFM) (wang2025fudoki; luo2025next; havasi2025edit; nguyen2025oneflow; deng2025uniform) studies on Large-Language Models (LLMs) or VLMs have demonstrated competitive performance on text reasoning and editing, as well as image and video generation, by enabling iterative token refinement and flexible generation trajectories guided by velocity field. These advances inspire us to transfer the same refinement principle to discrete action generation for robotic control.
We propose DFM-VLA, a VLA framework that unifies language, vision, and action in a discrete formulation, and employs a discrete action velocity field to enable holistic refinement of action sequences. As shown in Figure˜1, instead of treating action prediction as one-shot token generation, DFM-VLA models a token-level probability velocity field and performs iterative full-sequence refinement. This formulation allows the model to repeatedly revisit previously updated positions and selectively correct uncertain tokens as context becomes more informative. For velocity-field construction, we explore two variants, namely an auxiliary velocity head formulation that predicts velocities from model hidden states, and an action-embedding-guided formulation that defines semantically structured probability paths and derives kinetic-optimal velocities. For decoding, we adopt a two-stage strategy consisting of an iterative refinement stage for exploration and correction, followed by a deterministic validation stage for stable convergence.
We evaluate DFM-VLA on CALVIN, LIBERO, and real-world manipulation tasks. Across benchmarks, DFM-VLA consistently improves robot manipulation quality while maintaining strong inference efficiency. Ablation studies further indicate that the proposed design is robust to key scheduler settings, training data scale, and two-stage decoding allocation.
Our contributions are summarized as follows:
  • We identify the irreversible commitment problem in existing autoregressive and discrete diffusion VLA decoding, which limits token correction in robot manipulation.
  • We propose DFM-VLA, a discrete flow matching VLA framework that performs iterative full-sequence action refinement via token-level velocity.
  • We propose two strategies for velocity field construction, namely an auxiliary velocity head formulation and an action-embedding-guided formulation, and we systematically analyze these design variants.
  • We demonstrate strong empirical performance on CALVIN, LIBERO, and real-world tasks, with consistent gains in both action quality and inference efficiency.
Figure 2: Comparison of three discrete VLA paradigms. AR VLA uses causal attention to generate future action tokens from previously generated ones, $D$-VLA decodes only masked tokens, while DFM-VLA iteratively refines tokens over the full sequence.
2 Related Works
Autoregressive VLA.
Autoregressive VLA models(see Figure˜2) represent actions as discrete tokens, enabling LLM-style decoding and scalable training. Representative systems include RT-1 (brohan2022rt), OpenVLA (kim2024openvla), and LLaVA-VLA (song2026rethinking), which unify vision, language, and action prediction under the next-to-next objective. ReconVLA (song2025reconvla) improves perception by reconstructing the current frame to assist discrete action generation. Meanwhile, dynamic inference or efficiency tokenizer designs (e.g., Deer-VLA (yue2024deer) and FAST (pertsch2025fast)) target practical deployment constraints. Methods based on vector quantization (VQ) (wang2025vq; liu2025faster; dong2026actioncodec) provide a flexible alternative by learning discrete latent representations. UniVLA (univla) and WorldVLA (worldvla) discretize future frames to leverage dynamic visual information for action generation during both training and inference. These autoregressive models are effective but inherently sequential, which motivates our refinement alternative.
Discrete Diffusion VLA.
Discrete diffusion VLA models extend dLLM-style (yu2025discrete) parallel denoising to action sequences, enabling full-attention, multi-token decoding with iterative refinement. Representative methods differ in their corruption processes and efficiency strategies. PD-VLA (song2025accelerating) follows a BART-style noising scheme by substituting tokens with vocabulary items and learning to reconstruct the sequence, while Discrete Diffusion VLA (liang2025discrete) and LLADA-VLA (wen2025llada) adopt BERT-style masking with a dedicated mask token (bert). To reduce inference cost, CEED-VLA (song2025ceed) applies consistency distillation to shrink the number of denoising steps with minimal performance loss. On the data and modeling side, Dream-VLA (ye2025dream) scales discrete diffusion pretraining on the OXE (o2024open) datasets following OpenVLA (kim2024openvla). Likewise, UD-VLA (chen2025unified), dVLA (wen2025dvla), and MM-ACT (liang2025mm) incorporate visual or textual chain-of-thought (CoT)and jointly diffuse future frames, reasoning traces, and action tokens for unified perception–reasoning–action generation. However, these dVLA methods cannot perform token-level refinement, as their updates are restricted to masked positions rather than allowing full-sequence refinement.
Discrete Flow Matching LLM and VLM.
Discrete flow matching (gat2024discrete) provides a principled framework for modeling probability paths over discrete tokens, with theoretical work (shaul2024flow) establishing kinetic-optimal discrete paths and velocity formulations that generalize mask-based mixtures. Building on these foundations, several large-scale multimodal models adopt discrete flow matching for unified understanding and generation, including Fudoki (wang2025fudoki) and Next-Omni (luo2025next), demonstrating strong any-to-any or omnimodal capabilities. EditFlow (havasi2025edit) introduces explicit edit operations (insertion, replacement, deletion) with learned rates to enable flexible sequence refinement, while OneFlow (nguyen2025oneflow) extends this idea to concurrent mixed-modal and interleaved generation. URSA (deng2025uniform) further illustrates how discrete metric structures can guide token transitions in generative modeling. We introduce discrete flow matching to the action modality, leveraging iterative refinement via a velocity field to progressively enhance action generation.
3 Preliminary: Discrete Flow Matching
We briefly review the key concepts and notation of discrete flow matching (gat2024discrete; wang2025fudoki) that are used throughout the paper. DFM aims to transform a known source distribution $p(x)$ into a target data distribution $q(x)$ over a discrete space. We consider $x=(x^{1},x^{2},\ldots,x^{D})\in\mathcal{S}=\mathcal{T}^{D}$, where $D$ is the number of discrete variables and $\mathcal{T}=[K]=\{1,2,\ldots,K\}$ is the finite alphabet of possible values.
Probability Paths. Given a source distribution $p(x)$ and a target distribution $q(x)$ on a finite state space $\mathcal{S}$, DFM defines a family of time-indexed distributions $\{p_{t}(x)\}_{t\in[0,1]}$ that smoothly interpolates between $p$ and $q$, referred to as probability paths. Each $p_{t}(x)$ is constructed as $p_{t}(x)\coloneqq\sum_{x_{1}\in\mathcal{S}}p_{t}(x\mid x_{1})q(x_{1})$, where the conditional distribution factorizes across dimensions, i.e., $p_{t}(x\mid x_{1})\coloneqq\prod_{i=1}^{D}p_{t}(x^{i}\mid x_{1}^{i})$. Each term $p_{t}(x^{i}\mid x_{1}^{i})$ interpolates between the base distribution $p(x^{i})$ and a point mass $\delta_{x_{1}^{i}}(x^{i})$, i.e., $\delta_{x_{1}^{i}}(x^{i})=1$ if $x^{i}=x_{1}^{i}$ and 0 otherwise. A common choice is the mixture path used in (gat2024discrete; havasi2025edit), defined by a time-dependent scheduler $\kappa_{t}(x_{1}^{i})\in[0,1]$:
$p_{t}(x^{i}\mid x_{1}^{i})=(1-\kappa_{t}(x_{1}^{i}))p(x^{i})+\kappa_{t}(x_{1}^{i})\delta_{x_{1}^{i}}(x^{i}),$ (1)
where $\kappa_{0}(\cdot)=0$ and $\kappa_{1}(\cdot)=1$. The term $p_{t}(x^{i}\mid x_{1}^{i})$ is a conditional forward probability path that describes how the state $x^{i}$ evolves given $x_{1}^{i}$.
Probability velocities. To realize the prescribed probability path $p_{t}(x)$, we use a Continuous-Time Markov Chain (CTMC). Its dynamics are governed by a probability velocity $u_{t}$, also called the transition rate, which describes how the current state $x_{t}$ moves toward the target state $x_{1}$ over time. Each token $i$ is updated independently according to
$x_{t+h}^{i}\sim\delta_{x_{t}^{i}}(\cdot)+h\,u_{t}^{i}(\cdot\mid x_{t}^{i},x_{1}^{i}),$ (2)
where $u_{t}^{i}(\cdot\mid x_{t}^{i},x_{1}^{i})$ is the velocity field, a conditional rate function that governs the flow of probability from $x_{t}^{i}$ to $x_{1}^{i}$. Equation (2) can be viewed as a small perturbation of the point mass $\delta_{x_{t}^{i}}$ scaled by the step size $h$, which models discrete state transitions in continuous time. The velocity field is central to DFM, which defines the probability-path dynamics and is the main quantity learned during training.
4 Method
Figure 3: Overall architecture of DFM-VLA. Given language–vision context and noised action tokens $x_{t}$, the model predicts clean actions $x_{1}$ and learns the velocity field via $\mathcal{L}_{\text{ce}}$ or $L_{\text{head}}$}.
This section is organized into four parts. We first introduce the model architecture in Sec. 4.1. We then present two velocity-field constructions: 1) an auxiliary velocity head formulation in Sec. 4.2 and 2) an embedding-guided formulation in Sec. 4.3. Finally, we describe the two-stage inference procedure in Sec. 4.4.
4.1 Architecture
As shown in Figure˜3, DFM-VLA follows an architecture similar to UniVLA and adopts a unified discrete token formulation for language, vision, and action modalities. Specifically, we tokenize text using the Emu3 (wang2024emu3) tokenizer, and discretize visual observations with a VQ tokenizer (zheng2022movq) using a compression ratio of 4. Both third-person and wrist-view images are represented as $25\times 25=625$ tokens. For action encoding, we discretize continuous actions with FAST (pertsch2025fast) and further compress action tokens using Byte Pair Encoding (BPE), with an action vocabulary size of 1024. To explicitly distinguish modalities, we enclose image and action token spans with boundary markers, i.e., boi/eoi for begin/end of image and boa/eoa for begin/end of action. In this work, we denote the discretized instruction and observations as $l$, and apply noising and prediction only to the action modality.
4.2 Modeling Velocities by Additional Head
Drawing inspiration from EditFlow (havasi2025edit), which employs additional velocity heads to model three types of edit operations, namely insertion, replacement, and deletion, we explore an alternative formulation for velocity-field construction. We retain only the replacement operation, because the action token sequence length is predefined by our action chunk design and the three operations are fundamentally equivalent under appropriate reformulation (nguyen2025oneflow; havasi2025edit). Concretely, given noisy action tokens $x_{t}$ and context $l$, the backbone first produces hidden states, and the additional velocity head then maps these hidden states to replacement velocities:
$h_{t}=f_{\theta}(x_{t},l),\qquad u_{t}^{\theta}(\cdot\mid x_{t})=u_{t}^{\text{head}}(h_{t}),$ (3)
where $f_{\theta}$ denotes the backbone network and $u_{t}^{\text{head}}$ denotes the velocity prediction head.
$\small\mathcal{L_{\text{head}}}=\mathbb{E}_{t\sim\mathcal{U}[0,1],\,{x}_{1},{x}_{t}}\left[\sum_{x\neq x_{t}}u_{t}^{\theta}(x\mid x_{t})-\sum_{i=1}^{N}\mathbf{1}_{[x^{i}_{t}\neq x^{i}_{1}]}\,\log u_{t}^{\theta}\!\left(x^{i}\mid x^{i}_{t}\right)p_{1\mid t}\left({x}^{i}_{1}\mid{x}^{i}_{t},{l}\right)\right].$ (4)
Here, $\mathbf{1}_{[x_{t}^{i}\neq x_{1}^{i}]}$ is an indicator that equals 1 only when the current token differs from the target token, so the update loss is applied only to positions that still need refinement.
4.3 Action-Embedding-Guided Velocity Modeling
Building on recent advances in discrete flow matching (gat2024discrete; wang2025fudoki; deng2025uniform), we parameterize the probability path in a metric-induced form. Specifically, let $d:\mathcal{T}\times\mathcal{T}\to\mathbb{R}_{\geq 0}$ be a distance such that $d(x^{i},x_{1}^{i})=0$ if and only if $x^{i}=x_{1}^{i}$, where $d(\cdot,\cdot)$ is measured in the action token embedding space. We define the conditional path as
$p_{t}(x^{i}\mid x_{1}^{i})=\mathrm{softmax}\big(-\beta_{t}\cdot d(x^{i},x_{1}^{i})\big),$ (5)
where $\beta_{t}:[0,1]\to\mathbb{R}_{\geq 0}$ is a monotonic schedule with boundary conditions $\beta_{0}=0$ and $\beta_{1}=\infty$. We instantiate it as
$\beta_{t}=c\left(\dfrac{t}{1-t}\right)^{\alpha},\qquad t\in[0,1),$ (6)
where $c>0$ and $\alpha>0$ control how fast probability mass concentrates toward the target token over time. This formulation preserves semantic neighborhood structure such that tokens closer to $x_{1}^{i}$ receive larger probability as $t\to 1$.
Given this prescribed path, we adopt the kinetic-optimal velocity obtained by minimizing transport energy under the flow constraints (wang2025fudoki; deng2025uniform; luo2025next):
$u_{t}^{i}(x^{i},z\mid x_{1})=p_{t}(x^{i}\mid x_{1}^{i})\,\dot{\beta}_{t}\,[d(z^{i},x_{1}^{i})-d(x^{i},x_{1}^{i})]_{+}$ (7)
where $\cdot]_{+}=\max\{\cdot,0\}$, $z^{i}$ is token in the vocabulary $\mathcal{T}$, and $\dot{\beta}_{t}$ is the derivative of $\beta_{t}$ w.r.t. $t$. This velocity moves probability mass from $z^{i}$ to $x^{i}$ only when $x^{i}$ is closer to $x_{1}^{i}$ than $z^{i}$, yielding a monotonic refinement process toward the clean target token.
Training. Given language–vision context $l$ and corrupted action tokens $x_{t}$, the model predicts the target action sequence $x_{1}$ by outputting per-position categorical logits. We optimize the expected cross-entropy:
$\mathcal{L_{\text{ce}}}=\mathbb{E}_{t\sim\mathcal{U}[0,1],\,{x}_{1},{x}_{t}}\left[-\log p_{1\mid t}\left({x}_{1}\mid{x}_{t},{l}\right)\right].$ (8)
where $p_{1|t}^{\theta}(\cdot\mid x_{t},l)$ denotes the predicted categorical distribution of the model at each action token position.
As shown in Figure˜5, the embedding-guided objective ($\mathcal{L}_{\text{ce}}$) converges faster and achieves better final performance than the head-based objective ($\mathcal{L}_{\text{head}}$). Unless otherwise specified, all experiments use the embedding-guided setting.
4.4 Inference
Figure 4: Visualization of a single decoding step in the iterative refinement stage. After predicting final state $x^{\text{pred}}_{1}$, the model does not directly output final action tokens. Instead, it constructs a velocity field to compute transition rates and selectively updates tokens to next state $x_{t+h}$ at each step.
We perform inference in two stages: an iterative refinement stage followed by a deterministic validation stage, with $T_{\mathrm{fine}}$ and $T_{\mathrm{val}}$ decoding steps, respectively.
In the iterative refinement stage, we employ an Euler discretization of the continuous-time Markov chain (CTMC) process $(x_{t})_{0\leq t\leq 1}$, following the approach in(deng2025uniform). As illustrated in Figure˜4, for each coordinate $i$ and each step from $t$ to $t+h$, we perform:
  • Sample $x_{1}^{i}\sim p_{1|t}^{i}(\cdot\mid x_{t},l)$ from the model;
  • Compute the total outgoing rate $\lambda^{i}=\sum_{x^{i}\neq x_{t}^{i}}u_{t}^{i}(x^{i},x_{t}^{i}\mid x_{1}^{i})$ using Eq. 7 or Eq. 3;
  • Draw $Z^{i}_{\text{change}}\sim U[0,1]$;
  • Update $x_{t+h}^{i}$: if $Z^{i}_{\text{change}}\leq 1-e^{-h\lambda^{i}}$, sample from $\frac{u_{t}^{i}(\cdot,x_{t}^{i}\mid x_{1}^{i})}{\lambda^{i}}(1-\delta_{x_{t}^{i}}(\cdot))$; otherwise keep $x_{t+h}^{i}=x_{t}^{i}$.
Here, $\lambda^{i}$ is the total transition intensity out of the current token $x_{t}^{i}$, so the jump probability $1-e^{-h\lambda^{i}}$ increases with $\lambda^{i}$. When a jump occurs, normalized rates favor states with larger velocity flow, which typically move the token closer to the predicted clean state $x_{1}^{i}$. Repeating this process over time enables iterative correction across the full sequence. Compared with mask-based discrete diffusion methods (yedreamVLA), this decoding scheme allows previously updated tokens to be revised again in later iterations, rather than being permanently fixed after being predicted.
During the validation stage, we adopt a greedy decoding strategy to improve stability in the final refinement steps. In the final $T_{\mathrm{val}}$ decoding steps, we disable stochastic jumps and switch to greedy decoding,
$x_{t+h}=\arg\max\mathrm{softmax}(p_{1|t}(\cdot|x_{t})),$ (9)
This hybrid design preserves exploratory refinement in earlier iterations while enforcing deterministic convergence near the end, leading to more stable validation-time trajectories and improved reproducibility. The concise two-stage decoding procedure is summarized in Algorithm˜1.
Algorithm 1 Two-Stage Decoding of DFM-VLA
1:Input: predictor $p_{\theta}$, context $l$, steps $T_{\mathrm{fine}},T_{\mathrm{val}}$, action vocabulary $\mathcal{V}$
2: Sample $x_{0}\sim\mathrm{Uniform}(\mathcal{V})$; set $T\leftarrow T_{\mathrm{fine}}+T_{\mathrm{val}}$
3:for $k=1$ to $T$ do
4:  $t\leftarrow(k-1)/T$, $h\leftarrow 1/T$
5:  $\hat{x}_{1}\sim p_{\theta}(\cdot\mid x_{t},l)$
6:  if $k\leq T_{\mathrm{fine}}$ then
7:   Compute velocity $u_{t}$ from $\hat{x}_{1}$ (Eq. 7 or Eq. 3)
8:   Update $x_{t+h}$ by a CTMC Euler step
9:  else
10:   Update $x_{t+h}\leftarrow\arg\max p_{\theta}(\cdot\mid x_{t},l)$
11:  end if
12:end for
13:Output: action sequence $x_{1}$
[…]
Table 1: Comprehensive Evaluation of Long-Horizon Robotic Manipulation on the CALVIN Benchmark. UniVLA denotes the variant without historical frames for fair comparison.
Method Task Tasks Completed in a Row Avg. Len $\uparrow$
1 2 3 4 5
[…]
Table 2: Evaluation and comparison on the LIBERO benchmark.
Method Spatial Object Goal Long Average
Octo (octo_2023) 78.9% 85.7% 84.6% 51.1% 75.1%
SpatialVLA (qu2025spatialvla) 88.2% 89.9% 78.6% 55.5% 78.1%
CoT-VLA (zhao2025cot) 87.5% 91.6% 87.6% 69.0% 81.1%
WorldVLA (worldvla) 87.6% 96.2% 83.4% 60.0% 81.8%
ThinkAct (huang2025thinkact) 88.3% 91.4% 87.1% 70.9% 84.4%
$\pi_{0}$-FAST (pertsch2025fast) 96.4% 96.8% 88.6% 60.2% 85.5%
MolmoAct (lee2025molmoact) 87.0% 95.4% 87.6% 77.2% 86.6%
FlowVLA (zhong2025flowvla) 93.2% 95.0% 91.6% 72.6% 88.1%
DreamVLA (dreamvla25) 97.5% 94.0% 89.5% 89.5% 92.6%
DFM-VLA+Head 94.2% 96.4% 92.8% 90.4% 93.5%
DFM-VLA+Embed 96.8% 98.8% 94.4% 92.6% 95.7%
5 Experiments
We conduct comprehensive experiments to evaluate the effectiveness of DFM-VLA on both simulation benchmarks and real-world robotic manipulation tasks. Our experiments are designed to answer the following research questions:
(RQ1) How does DFM-VLA compare with recent state-of-the-art VLA methods on CALVIN and LIBERO benchmarks (see Sections˜5.3 and 5.2)?
(RQ2) What empirical insights can guide key design choices for DFM-VLA, including schedule hyperparameters $c$ and $\alpha$, two-stage decoding allocation, and velocity-field construction (see Section˜5.4)?
(RQ3) What additional insights can we gain from in-depth analysis of decoding behavior, execution quality, and efficiency trade-offs (see Section˜5.5)?
(RQ4) Can DFM-VLA generalize effectively to real-world manipulation tasks (see Figure˜6)?
5.1 Setup
We initialize our model from checkpoints pretrained on robotic video data (univla) and keep the remaining settings consistent with UniVLA. Unless otherwise specified, we use a learning rate of $1\times 10^{-4}$ and a batch size of 8. All training and inference are conducted on 8 NVIDIA H100 GPUs. For simulation benchmarks (CALVIN and LIBERO), we train for 20k–32k steps depending on the setting, while each real-world task is trained for 5k steps. Unless otherwise noted, we follow (luo2025next) and use $c=3$ and $\alpha=1$ as the default setting.
5.2 Benchmarks
[…]
LIBERO. LIBERO (liu2023libero) is a simulation benchmark designed to stress-test cross-task generalization along multiple dimensions. It is organized into four suites: Spatial, Object, Goal, and Long. Spatial focuses on spatial-relation understanding under varied layouts, Object emphasizes object-level generalization, Goal evaluates goal-conditioned manipulation under different target configurations, and Long tests long-horizon compositional reasoning over multi-stage behaviors. Compared with CALVIN, LIBERO places stronger emphasis on broad generalization across task families. We report per-suite success rates and the overall average. Each suite contains 10 tasks, with 50 evaluation rollouts per task.
5.3 Comparison with SOTA
We report the main results on CALVIN and LIBERO in Tables˜1 and 2. On CALVIN, both variants of DFM-VLA outperform most strong VLA baselines, and DFM-VLA+Embed achieves the best overall average length of 4.44. Compared with unified AR VLAs (e.g., UniVLA), DFM-VLA+Embed improves 3-step and 5-step completion (0.892 vs. 0.862 and 0.780 vs. 0.690), indicating stronger long horizon consistency. It also outperforms continuous diffusion VLA models across all reported metrics, showing that discrete action refinement is competitive with optimization in continuous action spaces. Moreover, compared with AR baselines that enhance perception, such as ReconVLA, our method delivers gains of 0.19, showing that focused action iterative refinement is more effective. Although DFM-VLA+Head is slightly weaker than DFM-VLA+Embed, it still surpasses other baselines, further supporting the robustness of velocity guided action regeneration in long-horizon tasks.
On LIBERO, our method shows clear gains across all suites. DFM-VLA+Embed reaches the best overall average success rate of 95.7%, surpassing reasoning-guided discrete VLAs that rely on text or spatial chain of thought cues, for example, ThinkAct, MolmoAct, and CoT-VLA. It also improves over DreamVLA, which follows continuous diffusion with future visual prediction, by +3.1 points, and over FlowVLA, which uses autoregressive future optical flow cues, by +7.6 points. In particular, it achieves the highest scores on Object at 98.8% and Long at 92.6%, while maintaining strong performance on Spatial and Goal tasks. These results suggest that directly refining actions with previously generated coarse action signals is more effective than relying on auxiliary modality chain of thought guidance, leading to better long-horizon execution and cross-task generalization.
Figure 5: Comparison of velocity field constructions across training steps. The Embedding-guided variant consistently outperforms the Velocity Head variant in both convergence speed and final performance, reaching a state-of-the-art success rate of 95.7% on LIBERO with only 20k training steps.
5.4 Ablation Studies
We conduct ablation studies to isolate the effects of key design and training choices, including schedule hyperparameters $c$ and $\alpha$, two-stage decoding step allocation, and velocity-field construction. Detailed quantitative results are summarized in Tables˜3, 4 and 5.
Table 3: Ablation study on $c$ and $\alpha$.
$c$ $\alpha$ CALVIN $\uparrow$ LIBERO (%) $\uparrow$
1 1 4.38 95.0
3 1 4.44 95.7
5 1 4.40 95.2
3 0.5 4.40 95.6
Effect of $c$ and $\alpha$.
In Eq. 6, $c$ controls the overall scale of $\beta_{t}$, while $\alpha$ controls the temporal curvature of the schedule, determining how quickly refinement concentrates as $t\!\to\!1$. Table˜3 shows that DFM-VLA is robust to moderate changes of both parameters, with $c=3,\alpha=1$ giving the best trade-off. Under this setting, DFM-VLA achieves the Avg. Len. of 4.44 on CALVIN and average success rate of 95.7% on Libero. Increasing $c$ to 5 or decreasing $\alpha$ to 0.5 causes only minor fluctuations, indicating stable refinement dynamics across a reasonably wide hyperparameter region.
Table 4: Ablation on two-stage decoding step allocation with fixed total steps ($T_{\mathrm{fine}}+T_{\mathrm{val}}=16$).
$T_{\mathrm{fine}}$ $T_{\mathrm{val}}$ CALVIN $\uparrow$ LIBERO (%) $\uparrow$
16 0 4.37 94.8
15 1 4.39 95.2
14 2 4.44 95.7
12 4 4.41 95.4
8 8 4.35 95.1
Effect of Two-Stage Decoding.
[…]
Comparison of Velocity Field Constructions.
To further analyze the effect of velocity-field design, we compare the embedding-guided formulation with the head-based formulation across training steps. As shown in Figure˜5, the embedding-guided variant converges faster in the early stages of training and consistently reaches better final task performance. This trend indicates that embedding guidance provides more informative and smooth optimization signals, leading to both improved data efficiency and a stronger final policy.
5.5 In-Depth Analysis
Effectiveness of Decoding Method.
[…]
Method Avg. Len. $\uparrow$ Speed $\uparrow$
AR 4.18 50.2
$D$ 4.32 62.1
DD + Adap. Cache 4.27 118.3
DFM 4.42 60.2
DFM + Adap. Cache 4.40 121.0
[…]
Data Fraction AR $D$ DFM (Ours)
10% 1.71 2.84 3.21
50% 3.01 3.88 4.03
100% 4.18 4.32 4.44
We keep the architecture unchanged and compare AR, DD, and our method, while also validating the effectiveness of Adaptive Cache in inference. Results in Table˜6 show that DFM-VLA achieves the best action quality among the compared decoding strategies. With Adaptive Cache enabled, both DD and DFM obtain clear speedups, and DFM+Adaptive Cache delivers the fastest decoding while maintaining strong performance. These results indicate that DFM provides a favorable quality–efficiency trade-off under matched decoding steps.
Effect of training data size.
[…]
5.6 Real World Experiment
Figure 6: Left: The real-world robotic system for experiments. Right: The representative task demonstrations.
Setup. As shown in Figure˜6, we conduct real-world experiments on a bimanual AgileX platform equipped with two robotic arms, each with six degrees of freedom and a parallel gripper. The system is instrumented with three RGB cameras: one fixed camera mounted at an elevated central viewpoint and two wrist-mounted cameras, one on each arm.
Task Setting. We design three representative manipulation tasks: (1) bimanual collaborative lifting of a pot (Pot Lift); (2) grasping elongated vegetables and placing them into the pot, where the object position and pose are varied (Place Veg. to Pot); and (3) grasping a block and placing it onto a plate with varying height (Place Block to Plate). For each task, we collect 100 trajectories for training. During evaluation, we run 40 trials per task and report both task success rate and completion speed.
Results. We compare DFM-VLAagainst representative methods from three action-generation paradigms for real-world manipulation: the autoregressive baseline $\pi_{0}$-FAST, the discrete diffusion baseline Dream-VLA, and the continuous diffusion baseline RDT (liu2024rdt). Table 7 shows that DFM-VLA achieves the highest success rate on each task. Specifically, our DFM-VLA reaches a success rate of 77.5% on Pot Lift, which requires precise bimanual coordination; 70.0% on Place Veg. to Pot, which tests generalization to object positions and poses; and 65.0% on Place Block to Plate, which emphasizes spatial precision under varying plate heights. Overall, the average success rate of DFM-VLA is 70.8%, which surpasses RDT at 60.0% by 10.8 percentage points and Dream-VLA at 54.2% by 16.6 percentage points.
This result demonstrates that traditional discrete baselines (i.e., $\pi_{0}$-FAST and Dream-VLA) are weaker than the continuous diffusion baseline RDT because wrongly decoded action tokens cannot be refined, leading to error accumulation. In contrast, our method uses refined action guided by velocities over discrete action tokens to correct errors in time. Meanwhile, its unified discrete input-output space better preserves semantic understanding from autoregressive VLM backbones (liang2025discrete; wen2025dvla), yielding the best overall performance. These results further show that action refinement with discrete flow matching remains highly competitive and robust in real-world settings.
Table 7: Real-world success rate comparison (%) on three tasks.
Method Pot Lift Place Veg. to Pot Place Block to Plate Average
$\pi_{0}$-FAST (pertsch2025fast) 50.0 42.5 35.0 42.5
Dream-VLA (yedreamVLA) 57.5 62.5 42.5 54.2
RDT (liu2024rdt) 65.0 62.5 52.5 60.0
DFM-VLA+Head 70.0 67.5 60.0 65.8
DFM-VLA+Embed 77.5 70.0 65.0 70.8
6 Conclusion
In this paper, we present DFM-VLA, a discrete flow-matching VLA framework that enables iterative full-sequence action refinement through token-level velocity modeling. Our DFM-VLA can revisit and correct previously generated action tokens, improving action consistency for robotic manipulation. We further study two complementary velocity-field constructions and integrate a practical two-stage decoder. Extensive experiments demonstrate that DFM-VLA consistently delivers strong action quality while maintaining high inference efficiency.